OpenStudy (anonymous):

Verify the identity. cos 4x + cos 2x = 2 - 2 sin2 2x - 2 sin2 x

OpenStudy (cwrw238):

cos 4x + cos 2x = 1 - 2 sin^2 2x + 1 - 2 sIn^2 x = 2 - 2 sin^2 2x - 2 sin^2 x